OpenStudy (anonymous):

At an auction, a lucky wealthy collector paid $7,000,000 for three paintings: a Monet, a Picasso, and a van Gogh. The Monet cost $800,000 more than the Picasso. The price of the van Gogh was $200,000 more than twice the price of the Monet. What was the price of each painting

OpenStudy (anonymous):

Let the price of Picasso = $ x Monet price = x + 800,000 van Gogh price = 2(x + 800,000) + 200,000 = 2x + 1,600,000 + 200,000 = 2x + 1,800,000 ATQ, x + x + 800,000 + 2x + 1,800,000 = 7,000,000 ==> 4x + 2,600,000 = 7,000,000 ==> 4x = 7,000,000 - 2,600,000 ==> 4x = 4,400,000 ==> x = 4,400,000/4 ==> x = 1,100,000 So, price of Picasso = $ 1,100,000 Price of Monet = 1,100,000 + 800,000 = $ 1,900,000 Price of van Gogh = 2(1,100,000) + 1,800,000 = $ 4,000,000

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

u r welcome....